# The dimension and Bose distance of some BCH codes of length $\frac{q^{m}-1}{\lambda}$

### v1

@@ III Some known results on $q$-cyclotomic cosets modulo $q^{m}-1$
Throughout the rest of the paper, we always assume that $m$ is a positive integer, $h=\lfloor\frac{m}{2}\rfloor$, and $n=\frac{q^{m}-1}{\lambda}$, where $\lambda$ divides q1q-1. This implies that $\lambda n=q^{m}-1$. Additional, for a positive real number $a$, we define $N(a)$ as the number of integers in the interval [1,a1][1,a-1] that are not divisible by $q$, i.e., $N(a)=\lfloor a-1\rfloor-\lfloor(a-1)/{q}\rfloor$.
In this section, we briefly review some known results on the sizes and coset leaders of the $q$-cyclotomic cosets modulo $\lambda n=q^{m}-1$. These foundational results will be essential for deriving the main contributions of this paper. To maintain consistency, we follow the notation and terminology used in [39], which we now introduce for completeness. Let $Z_{q}$ denote the set of all non-negative integers less than $q$. Each integer $a\in[0,\lambda n]$ can be uniquely represented by its $q$-adic expansion as $a=\sum\limits_{\ell=0}^{m-1}a_{\ell}q^{\ell},$ where $a_{\ell}\in Z_{q}$ for all $\ell=0,1,\dots,m-1$. Let $Z_{q}^{m}$ be the set of all length-$m$ sequences of integers in $Z_{q}$. For simplicity, denote by $\mathbf{0}_{m}$ the sequence in $Z_{q}^{m}$ whose elements are all zero. We define an order on $Z_{q}^{m}$ using lexicographic order. Specifically, for any two sequences $U=(u_{m-1},\ldots,u_{1},u_{0})$ and $W=(w_{m-1},\ldots,w_{1},w_{0})$ in $Z_{q}^{m}$,
@@
    $U$ and $W$ are said to be equal, denoted by $U=W$, if $u_{\ell}=w_{\ell}$ for $\ell=0,\ldots,m-1,$
@@
    $U$ is less than $W$, denoted by $U<W$, if either $u_{m-1}<w_{m-1}$ or there exists an integer $i\in[0,m-2]$ such that $u_{i}<v_{i}$ and $u_{\ell}=w_{\ell}$ for all $\ell=i+1,\ldots,m-1,$ and
@@
The map $V$ from the set of all the integers in $0,\lambda n]$ to $Z_{q}^{m}$ is defined as
@@
We define the following sets
[…]
Let $m$ and $k$ be two integers such that $m\geq 4$ and $m-2h\leq k\leq\lfloor{(2m-1)}/{3}\rfloor-h$. When $m$ is an odd integer, for each integer $i\in[-k+1,k]$, we define $\mathcal{A}_{k}(i)$ as the set of all integers $a\in[q^{h+k},q^{h+k+1})$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$ that satisfies:
@@
When $m$ is an even integer, for each $i\in[-k,k]$, we define $\mathcal{B}_{k}(i)$ as the set of all integers $a\in[q^{h+k},q^{h+k+1})$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$ that satisfies the condition in (5) and the following:
@@
Additionally, we define $t(a)=\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-h-i}$ and $\alpha(a)=\sum\limits_{\ell=0}^{k+i-1}a_{\ell}q^{\ell}$ for each integer $a\in\mathcal{A}_{k}(i)$, and define $t(a)=\sum\limits_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-h-i}$ and $\alpha(a)=\sum\limits_{\ell=0}^{k+i}a_{\ell}q^{\ell}$ for each integer $a\in\mathcal{B}_{k}(i)$. Then we make the following remarks.
@@
$a=t(a)q^{h+i}+\alpha(a),$
Since $\lambda\mid q-1$, it follows that for any integer $a\in\mathcal{A}_{k}(i)$,
@@
$\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}=t(a)\cdot q^{2i-1}\quad\hbox{for }i\in[1,k],$
@@
$\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}+\sum\limits_{\ell=h+i}^{h-i}a_{\ell}q^{\ell-(h-i+1)}$ $=$ $t(a)\cdot q^{2i-1}\quad\hbox{for }i\in[-k+1,0].$
Noticing that $0\leq\sum\limits_{\ell=h+i}^{h-i}a_{\ell}q^{\ell-(h-i+1)}<1$ for $i\in[-k+1,0]$, we can conclude that (10) is equivalent to
@@
Theorem 1.
[39, Theorem 2] Let $m$ and $k$ be two integers such that $m\geq 4$ and $m-2h\leq k\leq\lfloor(2m-1)/{3}\rfloor-h.$
  • If $m$ is odd, then
  • If $m$ is even, then
Lemma 2.
[39, Corollary 2] Let $m$ be an even integer and $k\in\left[0,\lfloor(2m-1)/{3}\rfloor-h\right]$ be an integer. Suppose that $a\in\left[q^{h+k},q^{h+k+1}\right)$ is an integer. Then $a\in\mathcal{H}$ if and only if $V(a)$ has the form
with $a_{0}>0$ and $a_{k}>0$.
Theorem 2.
[39, Theorem 1] When $m$ is an odd integer, for any integer $a\in[1,q^{m-\lfloor{m}/{3}\rfloor}),$
$|C_{\lambda n}(a)|=m.$
When $m$ is an even integer, for any integer $a\in[1,q^{m-\lfloor{m}/{3}\rfloor}),$
IV Auxillary Lemmas
@@
Let $k$ and $a\leq q$ be two positive integers. Then
@@
Let $k$ be an integer, and let $a\leq q$ be a positive integer. Then
@@
Let $k$ and $a\leq q$ be two positive integers. Then
@@
V The dimension of $\mathcal{C}_{(q,n,\delta)}$for $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$
Let $m\geq 4$ be an integer and $h=\lfloor\frac{m}{2}\rfloor$. Let $\delta$ be an integer with $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$, let $k_{\delta}$ be the integer such that $q^{h+k_{\delta}}\leq\lambda(\delta-1)<q^{h+k_{\delta}+1}$, and let $\sum\limits_{\ell=0}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$ be the $q$-adic expansion of $\lambda(\delta-1)$. If $k_{\delta}\geq m-2h$, let $s_{\delta}$ be the smallest integer in $m-2h-k_{\delta},k_{\delta}]$ such that $\delta_{h+s_{\delta}}>0$ and define $w_{\delta}=\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$.
@@
$f(\delta)=\begin{cases}0,&\hbox{if }\delta\leq\frac{q^{h+1}-1}{\lambda}+1,\\ {\left(\begin{aligned} &\frac{(q-1)^{2}}{\lambda}(k_{\delta}-1)q^{2k_{\delta}-3}+\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\mu(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor\\ &+\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right]\end{aligned}\right)}&\hbox{if }\delta>\frac{q^{h+1}-1}{\lambda}+1,\end{cases}$ (13)
@@
where $\widetilde{\mu}(\delta)=\min\left\{\sum\limits_{\ell=0}^{h-k_{\delta}-1}\delta_{\ell}q^{\ell},\sum\limits_{\ell=h-s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell-(h-s_{\delta})}\right\}.$ We define the function $\tau(\delta)$ for each integer $\delta\in[2,\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1]$ by
@@
Additionally, we define the function $g(\delta)$ for each integer $\delta\in[2,\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1]$ by
@@
Let $m$ and $\delta$ be integers with $m\geq 4$ and $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$.
@@
    $\mathrm{dim}(\mathcal{C}_{(q,n,\delta)})=n-m\left[N(\delta)-f(\delta)\right].$ (17)
@@
    $\mathrm{dim}(\mathcal{C}_{(q,n,\delta)})=n-m\left[N(\delta)-\widetilde{f}(\delta)\right]-\frac{m}{2}g(\delta).$ (18)
@@
Let $m$ and $\delta$ be integers with $m\geq 4$ and $\frac{q^{m-h}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$
@@
Let $m$ and $\delta$ be integers with $m\geq 4$ and $\frac{q^{m-h}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$
@@
Let $m$ and $k$ be two integers with $m\geq 4$ and $1\leq k\leq\lfloor(2m-1)/3\rfloor-h$.
@@
Let $m\geq 4$ be an even integer, and let $k$ be an integer with $0\leq k\leq\lfloor(2m-1)/3\rfloor-h]$.
@@
Let $m\geq 4$ be an even integer, and let $k$ be an integer with $0\leq k\leq\lfloor(2m-1)/3\rfloor-h]$.
@@
If $2\leq\delta\leq\frac{q^{h+1}-1}{\lambda}+1$, then $\lambda\leq\lambda(\delta-1)\leq q^{h+1}-1$. By applying [35, Theorem 2.3], we can obtain
@@
If $\frac{q^{h+1}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$, then $q^{h+1}\leq\lambda(\delta-1)<q^{\lfloor(2m-1)/3\rfloor+1}$. This implies $1\leq k_{\delta}\leq\lfloor(2m-1)/{3}\rfloor-h$. By applying Theorem 1, we can conclude from Assertion 3 that for any integer $k\in\left[1,\lfloor({2m-1})/{3}\rfloor-h\right],$
$\left|\left[q^{h+k},q^{h+k+1}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|$ $=$ $\sum\limits_{i=-k+1}^{k}\left|\mathcal{A}_{k}(i)\cap\mathcal{D}_{\lambda}\right|$
$=$ $\frac{1}{\lambda}\left[(k-1)q^{2k-3}(q-1)^{3}(q+1)+q^{2k-1}(q-1)^{2}\right].$
@@
$\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|$ $=$ $\sum\limits_{{k}=1}^{k_{\delta}-1}\left|\left[q^{h+{k}},q^{h+{k}+1}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|$
$=$ $\sum\limits_{{k}=1}^{k_{\delta}-1}\frac{(q-1)^{3}}{\lambda}(q+1)({k}-1)q^{2{k}-3}+\sum\limits_{{k}=1}^{k_{\delta}-1}\frac{q^{2{k}-1}(q-1)^{2}}{\lambda}.$
@@
$\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|$ $=\frac{(q-1)^{2}}{\lambda}(k_{\delta}-1)q^{2k_{\delta}-3}+\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\mu(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor$
$+\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right].$
@@
Next, we establish equation (17). Noticing that $\lfloor({2m-1})/{3}\rfloor+1\leq m-\lfloor{m}/{3}\rfloor$, we have $\lambda(\delta-1)<q^{m-\lfloor{m}/{3}\rfloor}$. Therefore, we can apply Theorem 2 and Lemma 1 to obtain
$|C_{n}(a)|=|C_{\lambda n}(\lambda a)|=m\quad\hbox{for any integer }a\in[1,\delta-1]$ (32)
@@
If $2\leq\delta\leq\frac{q^{h}-1}{\lambda}+1$, then $2\lambda\leq\lambda(\delta-1)\leq q^{h}-1$. By applying [35, Theorem 2.3], we can conclude that
@@
If $\frac{q^{h}-1}{\lambda}+1<\delta\leq\frac{q^{h+1}-1}{\lambda}+1,$ then we have $q^{h}\leq\lambda(\delta-1)\leq q^{h+1}-1$. Since $m$ is even, we have
  • $k_{\delta}=0$;
@@
By substituting $k_{\delta}$ and $\mathcal{T}_{0}(\delta)$ as above into equations (20) and (23), we obtain

### v2

@@ III Some known results on $q$-cyclotomic cosets modulo $q^{m}-1$
Throughout the rest of the paper, let $m$ be a positive integer, let $\lambda$ be a positive divisor of q1q-1, and set $h=\lfloor\frac{m}{2}\rfloor$ and $n=(q^{m}-1)/\lambda$. For a positive real number $a$, we denote by $N(a)$ the number of integers in the range [1,a1][1,a-1] that are not divisible by $q$. Equivalently, $N(a)=\lfloor a-1\rfloor-\lfloor(a-1)/{q}\rfloor.$
In this section, we briefly review some known results on $q$-cyclotomic cosets modulo $\lambda n=q^{m}-1$. These foundational results will be essential for deriving the main contributions of this paper. To maintain consistency, we follow the notation and terminology used in [40], which we now introduce for completeness.
Let $Z_{q}=\{0,1,2,\ldots,q-1\}$ be the set of nonnegative integers strictly less than $q$, and let $Z_{q}^{m}=\{(a_{m-1},\ldots,a_{0}):a_{i}\in Z_{q}\}$ be the set of all length-$m$ sequences of integers in $Z_{q}$. For simplicity, we denote by $\mathbf{0}_{m}$ the sequence in $Z_{q}^{m}$ whose elements are all zero. We define an order on $Z_{q}^{m}$ by lexicographic order. Specifically, for any two sequences $U=(u_{m-1},\ldots,u_{1},u_{0})$ and $W=(w_{m-1},\ldots,w_{1},w_{0})$ in $Z_{q}^{m}$, we define:
@@
    $U$ is equal to $W$, denoted by $U=W$, if $u_{\ell}=w_{\ell}$ for $\ell=0,\ldots,m-1.$
@@
    $U$ is less than $W$, denoted by $U<W$, if either $u_{m-1}<w_{m-1}$ or there exists an integer $i\in[0,m-2]$ such that $u_{i}<v_{i}$ and $u_{\ell}=w_{\ell}$ for all $\ell=i+1,\ldots,m-1.$
@@
Note that each integer $a\in[0,\lambda n]$ can be uniquely represented by its $q$-adic expansion as $a=\sum\limits_{\ell=0}^{m-1}a_{\ell}q^{\ell},$ where $a_{\ell}\in Z_{q}$ for all $\ell=0,1,\dots,m-1$. We define a map $V$ from the set of integers in $0,\lambda n]$ to $Z_{q}^{m}$ as
@@
Assume that $m\geq 4$. Let $k$ be an integer such that $m-2h\leq k\leq\lfloor{(2m-1)}/{3}\rfloor-h$. If $m$ is an odd integer, for each integer $i\in[-k+1,k]$, we define $\mathcal{A}_{k}(i)$ as the set of all integers $a\in[q^{h+k},q^{h+k+1})$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$ satisfying the following conditions:
@@
If $m$ is an even integer, for each integer $i\in[-k,k]$, we define $\mathcal{B}_{k}(i)$ as the set of all integers $a\in[q^{h+k},q^{h+k+1})$ with $q$-adic expansion $\sum\limits_{\ell=0}^{h+k}a_{\ell}q^{\ell}$ satisfying the condition in (7) and the following:
@@
With the above preparations, we now present several known results on $q$-cyclotomic cosets modulo $\lambda n=q^{m}-1$ as follows.
Lemma 2.
[36, Theorem 2.3] If $m$ is odd, then for any integer $\leq q^{h+1}, aa$ is the coset leader of $C_{\lambda n}(a)$ if and only if $q\nmid a$. If $m$ is even, then for any integer $\leq 2q^{h}, aa$ is the coset leader of $C_{\lambda n}(a)$ if and only if $q\nmid a$.
Clearly, if an integer $a$ is divisible by $q$, then $a/q\in C_{\lambda n}(a)$. Therefore, $q\nmid a$ is a necessary condition for $a$ to be a coset leader modulo $\lambda n=q^{m}-1$. The above lemma shows that this condition is also sufficient for sufficiently small $a$. However, for larger values of $a$, there exist integers $a$ that are neither divisible by $q$ nor coset leaders modulo $\lambda n=q^{m}-1$. We denote by $\mathcal{S}$ the set such integers, that is,
[…]
In addition, we define $\mathcal{H}$ to be the set of integers in $1,\lambda n-1]$ that are coset leaders modulo $\lambda n$ and whose corresponding $q$-cyclotomic cosets have size $m/2$, i.e.,
[…]
Lemma 3.
[40, Theorem 2] Let $m$ and $k$ be two integers such that $m\geq 4$ and $m-2h\leq k\leq\lfloor(2m-1)/{3}\rfloor-h.$
  • If $m$ is odd, then
  • If $m$ is even, then
In [40, Theorem 1], we determined the sizes of certain $q$-cyclotomic cosets modulo $q^{m}-1$. There is a typo in the even-$m$ case: the upper bound $q^{m-\lfloor m/4\rfloor}$ should be $q^{m-\lfloor m/3\rfloor}$. For convenience, we restate this result below with the corrected range.
Lemma 4.
[40, Theorem 1] If $m$ is an odd integer, then for any integer $a\in[1,q^{m-\lfloor{m}/{3}\rfloor}),$
$|C_{\lambda n}(a)|=m.$
If $m$ is an even integer, then for any integer $a\in[1,q^{m-\lfloor{m}/{3}\rfloor}),$
The following result characterizes the integers $a$ within certain ranges for which the $q$-cyclotomic coset $C_{\lambda n}(a)$ has size $m/2$ when $m$ is even.
Lemma 5.
[40, Corollary 2] Let $m$ be an even integer and $k\in\left[0,\lfloor(2m-1)/{3}\rfloor-h\right]$ be an integer. Suppose that $a\in\left[q^{h+k},q^{h+k+1}\right)$ is an integer. Then $a\in\mathcal{H}$ if and only if $V(a)$ has the form
with $a_{0}>0$ and $a_{k}>0$.
Although the above results only describe the sizes and coset leaders of $q$-cyclotomic cosets modulo $\lambda n=q^{m}-1$, by applying Lemma 1, they also yield information on $q$-cyclotomic cosets modulo $n=(q^{m}-1)/\lambda.$ These facts will be used later to determine the dimension and Bose distance of BCH codes of length $n=(q^{m}-1)/\lambda$.
To facilitate the use of these results in subsequent sections, we introduce additional notation and make several remarks concerning the definitions of the sets $\mathcal{A}_{k}(i)$ and $\mathcal{B}_{k}(i)$. For each integer $a\in\mathcal{A}_{k}(i)$, we define
$t(a)=\sum_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-h-i}\quad\text{and}\quad\alpha(a)=\sum_{\ell=0}^{k+i-1}a_{\ell}q^{\ell}.$
Similarly, for each integer $a\in\mathcal{B}_{k}(i)$, we define
$t(a)=\sum_{\ell=h+i}^{h+k}a_{\ell}q^{\ell-h-i}\quad\text{and}\quad\alpha(a)=\sum_{\ell=0}^{k+i}a_{\ell}q^{\ell}.$
We then make the following remarks.
@@
$a=t(a)q^{h+i}+\alpha(a).$
Since $\lambda$ is a divisor q1q-1, it follows that for any integer $a\in\mathcal{A}_{k}(i)$,
@@
$\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}=t(a)\cdot q^{2i-1}\quad\hbox{for all integers }i\in[1,k],$
@@
$\sum\limits_{\ell=h-i+1}^{h+k}a_{\ell}q^{\ell-(h-i+1)}+\sum\limits_{\ell=h+i}^{h-i}a_{\ell}q^{\ell-(h-i+1)}=t(a)\cdot q^{2i-1}\quad\hbox{for all integers }i\in[-k+1,0].$
Noticing that $0\leq\sum\limits_{\ell=h+i}^{h-i}a_{\ell}q^{\ell-(h-i+1)}<1$ for all integers $i\in[-k+1,0]$, we conclude that (12) is equivalent to
@@
IV Auxiliary Lemmas
@@
Let $k$ and $a$ be two positive integers with $a\leq q$. Then
@@
Let $k$ be an integer, and let $a$ be a positive integer with $a\leq q$. Then
@@
Let $k$ and $a$ be two positive integers with $a\leq q$. Then
@@
V The dimension of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$
Suppose that $m\geq 4$. Let $\delta$ be an integer satisfying $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$. Define $k_{\delta}=\lfloor\log_{q}\lambda(\delta-1)\rfloor-h$. This implies that $q^{h+k_{\delta}}\leq\lambda(\delta-1)<q^{h+k_{\delta}+1}$. Write the $q$-adic expansion of $\lambda(\delta-1)$ as $\lambda(\delta-1)=\sum\limits_{\ell=0}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$. If $k_{\delta}\geq m-2h$, let $s_{\delta}$ denote the smallest integer in $m-2h-k_{\delta},k_{\delta}]$ such that $\delta_{h+s_{\delta}}>0$ and let $w_{\delta}=\sum\limits_{\ell=h+s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell}$.
@@ V The dimension of narrow-sense BCH codes of length $(q^{m}-1)/\lambda$
$f(\delta)=\begin{cases}0&\hbox{if }\delta\leq\frac{q^{h+1}-1}{\lambda}+1,\\ {\left(\begin{aligned} &\frac{(q-1)^{2}}{\lambda}(k_{\delta}-1)q^{2k_{\delta}-3}+\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\mu(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor\\ &+\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right]\end{aligned}\right)}&\hbox{if }\delta>\frac{q^{h+1}-1}{\lambda}+1,\end{cases}$ (15)
@@
where $\widetilde{\mu}(\delta)=\min\left\{\sum\limits_{\ell=0}^{h-k_{\delta}-1}\delta_{\ell}q^{\ell},\sum\limits_{\ell=h-s_{\delta}}^{h+k_{\delta}}\delta_{\ell}q^{\ell-(h-s_{\delta})}\right\}.$ Define the function $\tau(\delta)$ for each integer $\delta\in\left[2,\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1\right]$ by
@@
Additionally, we define the function $g(\delta)$ for each integer $\delta\in\left[2,\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1\right]$ by
@@
Let $m\geq 4$ be an integer, let $\lambda$ be a positive divisor of q1q-1, and let $\delta$ be an integer such that $2\leq\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$ Set $n=(q^{m}-1)/\lambda$.
@@
    $\mathrm{dim}\bigl(\mathcal{C}_{(q,n,\delta)}\bigr)=n-m\bigl[N(\delta)-f(\delta)\bigr].$ (19)
@@
    $\mathrm{dim}\bigl(\mathcal{C}_{(q,n,\delta)}\bigr)=n-m\bigl[N(\delta)-\widetilde{f}(\delta)\bigr]-\frac{m}{2}g(\delta).$ (20)
    Here $N(\delta)=\delta-1-\lfloor(\delta-1)/q\rfloor$ denotes the number of integers in $1,\delta-1]$ not divisible by $q$, and $f(\delta)$, $\widetilde{f}(\delta)$ and $g(\delta)$ are defined in (15), (16) and (18) respectively.
The proof of Theorem 1 proceeds via Assertions 15. Readers may find the proofs for these assertions in Appendices IM.
@@
Suppose that $\frac{q^{m-h}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$
@@
Suppose that $\frac{q^{m-h}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1.$
@@
Let $k$ be an integer such that $1\leq k\leq\lfloor(2m-1)/3\rfloor-h$.
@@
Suppose that $m$ is even. Let $k$ be an integer such that $0\leq k\leq\lfloor(2m-1)/3\rfloor-h]$.
@@
Suppose that $m$ is even. Let $k$ be an integer such that $0\leq k\leq\lfloor(2m-1)/3\rfloor-h]$.
@@
If $2\leq\delta\leq\frac{q^{h+1}-1}{\lambda}+1$, then $\lambda\leq\lambda(\delta-1)\leq q^{h+1}-1$. By applying Lemma 2, we obtain
@@
If $\frac{q^{h+1}-1}{\lambda}+1<\delta\leq\frac{q^{\lfloor(2m-1)/3\rfloor+1}-1}{\lambda}+1$, then $q^{h+1}\leq\lambda(\delta-1)<q^{\lfloor(2m-1)/3\rfloor+1}$. This implies $1\leq k_{\delta}\leq\lfloor(2m-1)/{3}\rfloor-h$. By applying Lemma 3, we can conclude from Assertion 3 that for any integer $k\in\left[1,\lfloor({2m-1})/{3}\rfloor-h\right],$
$\begin{split}\left|\left[q^{h+k},q^{h+k+1}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|&=\sum\limits_{i=-k+1}^{k}\left|\mathcal{A}_{k}(i)\cap\mathcal{D}_{\lambda}\right|\\ &=\frac{1}{\lambda}\left[(k-1)q^{2k-3}(q-1)^{3}(q+1)+q^{2k-1}(q-1)^{2}\right].\end{split}$
@@
$\begin{split}\left|\left[q^{h+1},q^{h+k_{\delta}}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|&=\sum\limits_{{k}=1}^{k_{\delta}-1}\left|\left[q^{h+{k}},q^{h+{k}+1}\right)\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|\\ &=\sum\limits_{{k}=1}^{k_{\delta}-1}\frac{(q-1)^{3}}{\lambda}(q+1)({k}-1)q^{2{k}-3}+\sum\limits_{{k}=1}^{k_{\delta}-1}\frac{q^{2{k}-1}(q-1)^{2}}{\lambda}.\end{split}$
@@
$\begin{split}\left|\left[1,\lambda(\delta-1)\right]\cap\mathcal{S}\cap\mathcal{D}_{\lambda}\right|=&\frac{(q-1)^{2}}{\lambda}(k_{\delta}-1)q^{2k_{\delta}-3}+\lfloor\frac{\mu(\delta)+w_{\delta}}{\lambda}\rfloor-\lfloor\frac{\lfloor\mu(\delta)/q\rfloor+w_{\delta}}{\lambda}\rfloor\\ &+\sum\limits_{i=-k_{\delta}+1}^{k_{\delta}}\sum\limits_{t\in\mathcal{T}_{i}(\delta)}\left[\lfloor\frac{\lfloor tq^{2i-1}\rfloor+t}{\lambda}\rfloor-\lfloor\frac{\lfloor tq^{2i-2}\rfloor+t}{\lambda}\rfloor\right].\end{split}$
@@
Next, we establish equation (19). Noticing that $\lfloor({2m-1})/{3}\rfloor+1\leq m-\lfloor{m}/{3}\rfloor$, we have $\lambda(\delta-1)<q^{m-\lfloor{m}/{3}\rfloor}$. Therefore, we can apply Lemmas 1 and 4 to obtain
$|C_{n}(a)|=|C_{\lambda n}(\lambda a)|=m\quad\hbox{for all integers }a\in[1,\delta-1]$ (34)
@@
If $2\leq\delta\leq\frac{q^{h}-1}{\lambda}+1$, then $\lambda\leq\lambda(\delta-1)\leq q^{h}-1$. By applying Lemma 2, we conclude that
@@
If $\frac{q^{h}-1}{\lambda}+1<\delta\leq\frac{q^{h+1}-1}{\lambda}+1,$ then we have $q^{h}\leq\lambda(\delta-1)\leq q^{h+1}-1$. In this case, we have
  • $k_{\delta}=s_{\delta}=0$;
@@
By substituting $k_{\delta}$, $s_{\delta}$ and $\mathcal{T}_{0}(\delta)$ as above into equations (22) and (25), we obtain